# The existence and upper bound for stable photon spheres in static spherically symmetric black holes

### v1

@@ I Introduction
Owing to the significance in astrophysical observations, it is important to study the property of photon spheres. In a spherically symmetric black hole of mass $M$, Hod proved that for Einstein gravity coupled to matter satisfying the weak energy condition and negative trace energy condition, the innermost photon sphere radius $r_{\mathrm{ph,in}}$ and total mass $M$ satisfy $r_{\mathrm{ph,in}}\leq 3M$ Hod:2013jhd . By using the same energy condition, Ref. [14] proved an relationship between innermost photon sphere and its shadow radius: $r_{\mathrm{sh,in}}\geq\sqrt{3}r_{\mathrm{ph,in}}$ Cvetic:2016bxi . A lower bound $r_{\mathrm{ph,in}}\geq 2M$ was also conjectured by Hod Hod:2011cxc but counterexample was found by Ref. Cvetic:2016bxi . Subsequent studies have also explored the bounds of the innermost photon sphere in various spacetime backgrounds Hod:2020pim ; Peng:2020shc ; Hod:2023jmx ; Liu:2024odv . For observational purposes, it is often relevant to consider the photon sphere with the largest impact parameter, as this is the one that dominates the black hole shadow Lima:2021las . Typically, this corresponds to the outermost photon sphere. Hod’s proof does not apply to the outermost one when there are multiple photon spheres, which do exist in black holes satisfying the dominant energy condition Liu:2019rib . Recently, a series of universal inequalities about outermost photon sphere was proposed Lu:2019zxb ; Feng:2019zzn ; Yang:2019zcn .
@@ III Upper bound on the radii of the stable photon spheres of black holes
    The condition $T\leq 0$ is not a standard energy condition but is physically motivated. It is satisfied by many conventional fields (e.g., electromagnetic fields, dust, radiation and hairy black-hole configurations in Einstein-matter models Nunez:1996xv ), excluding exotic matter such as stiff matter which has $p=\rho$ Chavanis:2014lra .

### v2

@@ I Introduction
Owing to the significance in astrophysical observations, it is important to study the property of photon spheres. In a spherically symmetric black hole of mass $M$, Hod proved that for Einstein gravity coupled to matter satisfying the weak energy condition and negative trace energy condition, the innermost photon sphere radius $r_{\mathrm{ph,in}}$ and total mass $M$ satisfy $r_{\mathrm{ph,in}}\leq 3M$ Hod:2013jhd . By using the same energy condition, Ref. [14] proved an relationship between innermost photon sphere and its shadow radius: $r_{\mathrm{sh,in}}\geq\sqrt{3}r_{\mathrm{ph,in}}$ Cvetic:2016bxi . A lower bound $r_{\mathrm{ph,in}}\geq 2M$ was conjectured by Hod Hod:2011cxc but a counterexample was found by Ref. Cvetic:2016bxi . Subsequent studies have explored the bounds of the innermost photon sphere in various spacetime backgrounds Hod:2020pim ; Peng:2020shc ; Hod:2023jmx ; Liu:2024odv . Ref. DiFilippo:2024poc further investigates the effect of the self-gravity of photon spheres, resulting from the accumulation of photon mass, on the spacetime background structure of a black hole. For observational purposes, it is often relevant to consider the photon sphere with the largest impact parameter, as this is the one that dominates the black hole shadow Lima:2021las . Typically, this corresponds to the outermost photon sphere. Hod’s proof does not apply to the outermost one when there are multiple photon spheres, which do exist in black holes satisfying the dominant energy condition Liu:2019rib . Recently, a series of universal inequalities about outermost photon sphere was proposed Lu:2019zxb ; Feng:2019zzn ; Yang:2019zcn .
@@ III Upper bound on the radii of the stable photon spheres of black holes
    The condition $T\leq 0$, which is consistent with the subdominant trace energy condition Bekenstein:2013ztp , is satisfied by many regular matter fields, such as electromagnetic fields, dust, radiation, and hairy black-hole configurations in Einstein-matter models Nunez:1996xv . However, it excludes certain extreme matter, such as stiff matter, which has $p=\rho$ Chavanis:2014lra .
@@ IV Discussion and conclusion
Acknowledgement
This work is supported by the Research Start-up Funding of Chengdu University of Technology (Grant No. 10912-KYQD2022-09307).